The following information pertains to JAE Corp. at January 1, Year 1: Common stock, $11 par, 13,000 shares authorized, 2,600b. How many shares of common stock had been issued at the end of the period? Issued shares at the end of the periodJournal entry worksheet 2 1 Issued 1,100 shares of $11 par common stock for $26 per share. Note: Enter debits before credits.Cash Common Stock |Beg. Bal Beg. Bal End. Bal End. Bal Treasury Stock PIC in Excess of Par, CS Beg. Bal Beg. Ba End. Bal End.

Sorry for the long post, just need some help :)
Answer #1

Outstanding shares at the end of the period = 2,600+1, 100-180+50 =3,570 b Issued shares = 2,600+1,100 = 3,700 C Transaction

Cash Beg. Bal. Beg. Bal. 1 Common stock 28,600 12,100 28,600 1,200 25,660 End. Bal. End. Bal. 40,700 Treasury stock PIC in ex

40,700 33,400 50 Stockholders equity Common stock, $11 par Paid in capital in excess of par, Common stock Paid in capital in